What is the ratio of lactic acid (Ka = 1.37x10-4) to lactate in a solution with pH =4.07?

Part 1:

[NH3]=

Part 2:

[NH4+]
Answer #1

Ka = 1.37*10^-4

pKa = - log (Ka)

= - log(1.37*10^-4)

= 3.863

use:

pH = pKa + log {[conjugate base]/[acid]}

4.07 = 3.863+log {[A-]/[HA]}

log {[A-]/[HA]} = 0.2067

[A-]/[HA] = 1.61

[HA]/[A-] = 0.621

Answer: 0.621